Please help me solve this equation: An apple pie and a pumpkin pie are being served at thanksgiving dinner. The apple pie is sliced first. One person eats 1/12 of the pie, the second person eats 1/24 of the pie, and a third person eats 1/12 of the pie. The pattern continues in this way. When the pumpkin pie is sliced for the first time, the first person eats 1/9 of the pie, the second person eats 1/18 of the pie, the third person eats 1/9 of the pie the pattern continues in this way. If both pies have been eaten, how many more people have eaten apple pie than pumpkin pie?
-thanks!!

Answer by stanbon(75887)

An apple pie and a pumpkin pie are being served at thanksgiving dinner. The apple pie is sliced first. One person eats 1/12 of the pie, the second person eats 1/24 of the pie, and a third person eats 1/12 of the pie. The pattern continues in this way.
------
(1/12)+ (1/24) + (1/12) + (1/24) = 3/12 =1/4
4 persons eat 1/4 of the pie
1 pie = 4*4 = 16 persons
=================================
When the pumpkin pie is sliced for the first time, the first person eats 1/9 of the pie, the second person eats 1/18 of the pie, the third person eats 1/9 of the pie the pattern continues in this way. (1/9) + (1/18) + (1/9) + (1/18) = 3/9 = 1/3 pie
4 persons eat 1/3 pie
1 pie = 3*4 = 12 persons
-------------------------------------
If both pies have been eaten, how many more people have eaten apple pie than pumpkin pie?
Ans: 16-12 = 4
--------------------
Cheers,
Stan H.
================

------------------------------------